From GlueXWiki
Jump to: navigation, search

September 1, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • Package configurations
    • First package tests (Lubomir)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126 (cell#2) (Beni, Lubomir)
  5. Other


Participants: Bill, Eugene, Dave, Chris, Simon, Beni, Glenn, and Lubomir.


- During Irene the power was always on and the humidity and particle counts were OK on Monday.

- Dave: wire frame #6 is out of the stringing table and now working on #7: it was strung yesterday by Casey and Al in ~4 hours (record). Then the wires were taped and now doing position measurements. Aanatoly finished replacing the caps on frame #7. Cathodes: Tina and Mike tensioned type-1 cathode from set#7 and put the daughter cards on type-2 from set#6. Anatoly finished testing all the 80 foil sets. Wire frame #6 will be populated by Chris tomorrow and then (possibly on Monday) the central wires will be deadened.

- In the construction tracking linked above, at the second page, there's a table about the wire frames explaining differences in the material/procedures used: Epon or Epo-lite, with or without Humi-seal, capacitors replaced/not replaced, and whether the wires were deadened. Starting from frame #6 and up (6 is the consecutive number, not the frame number that is labeled) on all the frames the caps were/will be replaced without Humi-seal and Epon was/will be used instead of Epo-lite. If we exclude the first four wire planes as spares, there's only one (#5) production frame on which the caps were not replaced, but still Epon was used. We discussed if we want to replace the caps on all the frames and how, or if we just want to build new wire frames using the latest technology. Bill: we need to replace the old caps everywhere. However it will be very difficult, time consuming and risky: one can easily damage wires especially on the frames with Humi-seal on the caps. Eugene: it might be better just to make new wire frames, but it has to be evaluated. Lubomir will collect more information for the next meeting what needs to be purchased to produce extra wire frames.

- We started stacking the first production package. The first cell was installed yesterday and since then is being flushed with gas. The current between sense and field wires dropped by a factor of 10 in ~12 hours. It is explained by the fact that we used Epon; this is to be compared with more than one week for cell#2 when we used Epolite. In the moment the currents are of the same order (~700nA) as those that we have on cell#2 in 126. Before stringing we had 50-60nA on the capacitors only.

- The iron-55 source to be used for testing at Blue Crab was shipped yesterday (a week delay).


- Flatness measurements were done on the first package and the results are attached above. The aluminum fixture was measured with a height gauge lifted with two precision blocks on the granite table. Then the flatness of the gusset ring was measured in the same way; the ring was sitting on the hubs and it was tightened also on the top with nuts. The overall variations of the height of the gusset ring is +/-6 mils or +/- 150 microns. This is so far acceptable but of course important is the flatness when the whole package is stacked.


- HV cap problem: Chris contacted the procurement and sent a complaint letter with pictures to the company. The question is if we can ask the company about some sort of compensation if their product doesn't meet the technical requirements.

Chamber testing at 126 (cell #2)

- Lubomir showed drift time spectra (FDC logbook page 576) for two very high HV for this gas mixture (pre-mixed 89.8/10.2 Ar/CO2): 1800V and 1875V. Certainly there's strong dependence of the drift time on the HV and at 1800V it's much flatter. It will be interesting to compare with the results at 1650V that Beni has. Also, additional studies of the deadened area were done placing the 55Fe source in the middle of the chamber and triggering with one of the central wires. Plotted on page 577 are the number of counts on the top strips; note that 12 strips cover the part of the central wire that is deadened. One concludes that the counting on the strips at the edge of the area is reduced by a factor of 100. Note also that the source was 1cm above the wires and the gamma flux in the middle is ~9 times higher than at the edge of the deadened area. At the same time the strips in the middle don't reach the central wire by 1.8 cm.

- Beni: one of the two pressure sensors was fixed (solder ball problem) and the other one was sent also back to the company for fixing. Will need 5-6 like these but not sure whether to get them from the same company.


- Sarting September 16 we will have a Russian graduate student from MePhi (Moscow), Vladimir Berdnikov, working with us on the FDC.